A small artery has a length of 1.1x10 m and a radius
of 2.5x 10 m . If the pressure drop across the artery is
1.3.kPa, what is the flow rate through the artery? (Assume
that the temperature is 37°C.)
1
Expert's answer
2019-11-18T10:43:30-0500
"\eta = 2.1\cdot 10^{-3} \frac{Ns}{m^2}"

Hagen–Poiseuille equation:

"\Delta p =\frac{ 8\eta LQ}{\pi R^4}"

"(1.3\cdot 10^{3}) =\frac{ 8(2.1\cdot 10^{-3})(1.1\cdot 10^{-3})Q}{\pi (2.5\cdot 10^{-5})^4}"

"Q=8.7\cdot 10^{-11}\frac{m^3}{s}"